### v1

# Avoiding the Price of Adaptivity: Inference in Linear Contextual Bandits via Stability

Abstract
Statistical inference in contextual bandits is complicated by the adaptive, non-i.i.d. nature of the data. A growing body of work has shown that classical least-squares inference may fail under adaptive sampling, and that constructing valid confidence intervals for linear functionals of the model parameter typically requires paying an unavoidable inflation of order $\sqrt{d\log T}$. This phenomenon—often referred to as the price of adaptivity—highlights the inherent difficulty of reliable inference under general contextual bandit policies.
A key structural property that circumvents this limitation is the stability condition of Lai and Wei (lai1982least), which requires the empirical feature covariance to concentrate around a deterministic limit. When stability holds, the ordinary least-squares estimator satisfies a central limit theorem, and classical Wald-type confidence intervals—designed for i.i.d. data— become asymptotically valid even under adaptation, without incurring the $\sqrt{d\log T}$ price of adaptivity.
In this paper, we propose and analyze a penalized EXP4 algorithm for linear contextual bandits. Our first main result shows that this procedure satisfies the Lai–Wei stability condition and therefore admits valid Wald-type confidence intervals for linear functionals. Our second result establishes that the same algorithm achieves regret guarantees that are minimax optimal up to logarithmic factors, demonstrating that stability and statistical efficiency can coexist within a single contextual bandit method. Finally, we complement our theory with simulations illustrating the empirical normality of the resulting estimators and the sharpness of the corresponding confidence intervals.
1 Introduction
Statistical inference under adaptive data collection has become a central challenge in modern learning systems. Unlike classical settings where data are sampled independently of past observations, adaptive algorithms—such as contextual bandits and reinforcement learning policies—select actions based on previously observed losses. This feedback loop creates intricate dependencies that distort the distribution of both covariates and errors, often rendering standard inferential tools invalid. Even when estimators remain consistent, their asymptotic distributions can deviate substantially from classical theory, complicating uncertainty quantification and hypothesis testing (dickey1979distribution; lai1982least; zhang2020inference; deshpande2023online; khamaru2021near).
A recurring theme in recent work is that valid inference is possible using when the data-collection rule exhibits sufficient regularity or stability. At a high level, a bandit algorithm is stable if the long-run behavior of the algorithm settles into a predictable pattern, despite ongoing adaptation. When such structure exists, it becomes possible to characterize limiting distributions of estimators, derive confidence sets, and recover analogs of the classical central limit theorem (lai1982least). Recent works have demonstrated that this property is satisfied for the Upper Confidence Bound-type of algorithms (kalvit2021closer; fan2022typical; khamaru2024inference; han2024ucb), and a variant of the Thompson Sampling (halder2025stable; fan2022typical). Recent work of fan2025statistical show that the LinUCB (li2010contextual; abbasi2011improved) algorithm algorithm also satisfy this stability condition for linear contextual bandit problem.
In this paper, we investigate stability properties of bandit algorithms for a linear contextual bandit problem. Formally, at each round $t$, the learner observes a context $x_{t}\in\mathcal{X}$ and selects an action $a_{t}\in\mathcal{A}$ based on past data and context $x_{t}$. Concretely, let $\mathcal{F}_{t-1}:=\sigma\left(x_{1},a_{1},\ell_{1},\ldots,x_{t-1},a_{t-1},\ell_{t-1}\right)$ denote the sigma-field generated by the observations up to time t1t-1, then $a_{t}$ depends on $\mathcal{F}_{t-1}$ and $x_{t}$. Upon selecting an action $a_{t}$ we incur a loss according to a linear model:
$\ell_{t}\;=\;\langle\phi(a_{t},x_{t}),\,\beta^{\star}\rangle+\varepsilon_{t},$ (1)
where $\phi:\mathcal{A}\times\mathcal{X}\to\mathbb{R}^{d}$ is a known feature map and $\beta^{\star}\in\mathbb{R}^{d}$ is an unknown parameter to be learned. We assume that the noise sequence $\{\varepsilon_{t}\}_{t\geq 1}$ satisfies $\mathbb{E}[\varepsilon_{t}\mid\mathcal{F}_{t-1},x_{t}]=0$. Our goal is to construct confidence intervals for linear functional of the form $a^{\top}\beta^{\star}$, for any fixed vector $a\in^{d}$.
1.1 Price of Adaptivity:
Before we dive into more details, it is useful to compare the available method for constructing confidence intervals for $a^{\top}\beta^{\star}$. One popular approach of constructing confidence intervals for $a^{\top}\beta^{\star}$ is to use martingale concentration inequalities (de2004self; de2009self). Formally, let $\widehat{\beta}_{\lambda,T}$ denote the ridge-estimator with regularizer $\lambda$ based on data up to time $T$. Let
$\mathbf{V}_{T}\;=\;\lambda I_{d}+\sum_{t=1}^{T}\phi(a_{t},x_{t})\,\phi(a_{t},x_{t})^{\top}$
be the regularized design matrix. Given a target confidence $\alpha\in(0,1)$, a widely used approach, originating from the work of abbasi2011improved, is to construct a confidence interval
$\mathcal{I}^{\mathrm{APS}}_{T}(a)\;:=\;\Bigl[a^{\top}\hat{\widehat{\beta}}_{\lambda,T}\;\pm\;\mathrm{R}_{T}\,\sqrt{a^{\top}\mathbf{V}_{T}^{-1}a}\Bigr].$ (2)
Assuming $\|\beta^{\star}\|_{2}\leq S$, $\|\phi(x,a)\|_{2}\leq L$ and noise $\epsilon_{t}$ is $1$ sub-Gaussian, the factor $\mathrm{R}_{T}$ takes the following form
$\mathrm{R}_{T}\;=\;\sqrt{d\log\!\Bigl(\tfrac{TL}{\lambda}\Bigr)+\log(1/\alpha)}\;+\;\sqrt{\lambda}\,S.$ (3)
It is useful to compare the confidence interval (2) with a Wald-type interval, which is asymptotically exact when the data $(x_{t},a_{t})\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{P}$:
$\mathcal{I}^{\mathrm{Wald}}_{T}(a)\;:=\;\Bigl[a^{\top}\widehat{\beta}_{\mathrm{OLS}}\;\pm\;z_{1-\alpha/2}\,\widehat{\sigma}\,\sqrt{a^{\top}S_{T}^{-1}a}\Bigr],$ (4)
[…]
$\frac{\mathrm{width}\bigl(\mathcal{I}^{\mathrm{APS}}_{T}(a)\bigr)}{\mathrm{width}\bigl(\mathcal{I}^{\mathrm{Wald}}_{T}(a)\bigr)}\;\approx\;\sqrt{d\log T},$ (5)
Put simply, for contextual bandit problems the confidence intervals are $\sqrt{d\log T}$ times wider.
It is natural to ask whether one might construct any other confidence interval based on some other estimator. A recent line of work by lattimore2023lower; khamaru2021near; vakili2021open show that this enlagrement in the confidence interval length by a factor of $\sqrt{d\log T}$ is necessary when $d\geq 2$. State differently, this enlargement in confidence interval for contextual bandit is unavoidable in a worst case sense.
1.2 Validity of Wald’s Interval via Stability:
The comparison above highlights the price of adaptivity: without further structure, any confidence interval for $a^{\top}\beta^{\star}$ must inflate by a factor of order $\sqrt{d\log T}$, reflecting the worst–case distortions introduced by adaptive sampling. A natural question is whether this inflation is intrinsic to all adaptive procedures, or whether additional regularity in the data–collection rule can restore the validity of classical Wald’s confidence interval (4).
Following the seminal work of lai1982least, this regularity can be formalized through the notion of stability.
Definition 1.
We call a bandit algorithm $\mathcal{A}$ stable if there exists a sequence of non-random positive definite matrix $\{{\Sigma^{\star}_{T}}\}$ such that
${\Sigma^{\star}_{T}}^{-1}\mathrm{S}_{T}\;\stackrel{{\scriptstyle p}}{{\rightarrow}}\;\mathbf{I}$ (6)
Theorem 3 of lai1982least ensures that if a bandit algorithm $\mathcal{A}$ is stable, then the least square estimator $\widehat{\beta}_{\mathrm{OLS}}$ is asymptotically normal:
$\frac{1}{\widehat{\sigma}}\cdot\mathrm{S}_{T}^{-\frac{1}{2}}\left(\widehat{\beta}_{\mathrm{OLS}}-\beta^{\star}\right)\xrightarrow{d}\mathcal{N}\left(0,\mathbf{I}\right)$ (7)
where $\widehat{\sigma}$ is a consistent estimator  [^2] For instance, one might use the estimator from (lai1982least, Lemma 3). for noise standard deviation $\sigma$. An immediate consequence of the above using Slutsky’s theorem is that
$\lim_{T\rightarrow\infty}\mathbb{P}\left(\mathcal{I}^{\mathrm{Wald}}_{T}(a)\ni a^{\top}\beta^{\star}\right)=1-\alpha.$
In other words, Wald’s confidence interval for $a^{\top}\beta^{\star}$ is valid as long as the underlying bandit algorithm $\mathcal{A}$ is stable, and we do not pay the $\sqrt{d\log T}$ price of adaptivity. Our goal in this paper is to propose and analyze an algorithm which is simultaneously stable and provide good regret.
1.3 Contributions
In this paper, we study the problem of constructing confidence intervals for linear functionals of the model parameters in a linear contextual bandit problem. Section 2 provides a detailed description of the contextual bandit framework we consider, while Section 3 introduces the algorithm we analyze. The procedure, stated formally in Algorithm 1, is a penalized variant of the classical EXP4 algorithm designed to enforce stability while allowing adaptive exploration. Our first main result, stated in Theorem 1, establishes that the penalized EXP4 algorithm is stable in the sense of lai1982least (see definition (1)), enabling the construction of asymptotically valid Wald-type confidence intervals. Our second result, detailed in Theorem 2 , analyzes the regret of the proposed algorithm. We show that, despite being explicitly designed for inferential stability, the procedure simultaneously achieves minimax–optimal regret guarantees up to logarithmic factors. Thus, stability and statistical efficiency can coexist within a single contextual bandit algorithm. Section 6 contains proofs of our main theorems with some technical details deferred to the Appendix. Finally, in Section 5, we complement our theoretical findings with numerical experiments demonstrating the empirical validity of the Wald confidence intervals produced by Algorithm 1.
1.4 Related Work
The challenges of performing statistical inference under adaptive data collection have been well documented in recent years. Because adaptive policies bias the distribution of covariates and errors, classical inferential procedures may become invalid. This breakdown has been observed empirically (xu2013estimation; villar2015multi) and supported by theoretical analyses (nie2018adaptively; shin2019sample; shin2021bias). To counteract these issues, several methodological approaches have been proposed, including online debiasing techniques (khamaru2021near; chen2022debiasing; kim2023double) and procedures based on inverse propensity scores (hadad2021confidence; deshpande2018accurate; zhang2022statistical; nair2023randomization; leiner2025adaptive). These methods address inference in adaptive environments more broadly, but do not directly resolve the structural constraints that give rise to the $\sqrt{d\log T}$ “price of adaptivity” in contextual bandits.
A distinct line of work focuses on identifying conditions under which classical, Wald-type asymptotic inference is restored despite adaptivity. The seminal paper of lai1982least introduced the notion of stability, showing that if the empirical design covariance converges to a deterministic limit, then least-squares estimators satisfy a central limit theorem and Wald confidence intervals regain asymptotic validity. This perspective has motivated a growing literature on stability-based inference for bandit algorithms (kalvit2021closer; khamaru2024inference; fan2022typical; fan2024precise; han2024ucb; halder2025stable; fan2025statistical). However, subsequent work has shown that many commonly used bandit algorithms fail to satisfy the Lai–Wei stability condition, leading to substantial under-coverage when Wald intervals are applied naively (fan2024precise; praharaj2025instability). This underscores the central question motivating our work: can one design adaptive algorithms that are simultaneously stable and statistically efficient?
Contextual and adversarial bandit algorithms are frequently derived from mirror descent (lattimore2020bandit), a framework that has been applied extensively in online optimization (abernethy2009competing; audibert2014regret; bubeck2018sparsity; wei2018more) and that underlies widely used adversarial bandit methods such as EXP3 (auer1995gambling), EXP4 (auer2002nonstochastic), Tsallis-INF (zimmert2021tsallis; masoudian2021improved), and OFTRL (pmlr-v134-ito21a). Despite their strong regret guarantees, the stability properties of these algorithms are not well understood, and existing analyses suggest that many of them may not support valid Wald-type inference.
Finally, an alternative approach to inference in adaptive settings relies on non-asymptotic, anytime-valid confidence intervals constructed via concentration inequalities for self-normalized martingales. This line of work builds on the foundational results of de la Peña et al. (de2004self; de2009self) and includes several refined analyses (abbasi2011improved; howard2020time; waudby2024anytime). These intervals hold uniformly over time and do not rely on asymptotic arguments, but the price paid is typically much wider confidence intervals compared to those achievable under stability, reflecting the worst-case nature of anytime-valid guarantees.
1.5 Notation
For any matrix $A$, let $\|A\|_{\mathrm{op}}$ and $\|A\|_{F}$ denote the operator norm and the Frobenius norm, respectively. For any vector $v\in\mathbb{R}^{d}$, the standard $\ell_{p}$-norm for $p\in(0,\infty)$ is denoted by $\|v\|_{p}$. For a real-valued random variable $u$, we define $\|u\|_{p}:=\mathbb{E}\!\left[|u|^{p}\right]^{1/p}.$ Given a fixed weight vector $w\in\mathbb{R}^{d}$, we define a weighted norm on $\mathbb{R}^{d}$, denoted by $\|\cdot\|_{w,\ast}$, as $\|v\|_{w,\ast}^{2}:=\sum_{i=1}^{d}w_{i}v_{i}^{2},$ where $v\in\mathbb{R}^{d}.$ For two nonnegative sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $b_{n}\gg a_{n}$ if $\frac{b_{n}}{a_{n}}\to\infty\quad\text{as }n\to\infty.$ Throughout, we suppress absolute constants in inequalities by using the notation $\lesssim$ and $\gtrsim$. The Loewner partial order on symmetric matrices is denoted by $\preceq$.
2 Problem Setup
We consider a linear contextual bandit problem with a finite action set $\mathcal{A}$. At each round $t=1,\ldots,T$, the learner observes a context vector $x_{t}\in\mathcal{X}$, drawn i.i.d. from a distribution $\mathcal{P_{X}}$, selects an action $a_{t}\in\mathcal{A}$, and receives a random loss
$\ell_{t}\;=\;\langle\beta^{\star},\,c(x_{t},a_{t})\rangle+\varepsilon_{t},$ (8)
where $c(x_{t},a_{t})\in\mathbb{R}^{d}$ is a known feature representation of the pair $(x_{t},a_{t})$, and $\beta^{\star}\in\mathbb{R}^{d}$ is an unknown parameter vector. We use $\mathcal{F}_{t}:=\sigma(x_{1},a_{1},\ell_{1},\ldots,x_{t},a_{t},\ell_{t})$ to denote the $\sigma$-field generated by observations up to time $t$. We assume that the noise sequence $\{\varepsilon_{t}\}_{t=1}^{T}$ satisfies
$\mathbb{E}[\varepsilon_{t}\mid\mathcal{F}_{t-1},x_{t}]=0.$ (9)
At each round, the learner chooses an action by sampling from a mixture of $K$ base experts $\{\pi_{1},\ldots,\pi_{K}\}$, where each $\pi_{k}(\cdot\mid x_{t})$ defines an arbitrary (possibly stochastic) policy mapping the context $x_{t}$ to a distribution over actions. The learner maintains mixture weights $w_{t}=(w_{t,1},\ldots,w_{t,K})$ belonging to the $\varepsilon$-simplex
$\Delta_{\varepsilon}=\left\{\,w\in\mathbb{R}^{K}_{\geq 0}:\sum_{k=1}^{K}w_{k}=1,\;w_{k}\geq\varepsilon\,\right\}.$ (10)
We take $\epsilon$ to be a small positive tuning parameter that decays with the number of rounds $T$. Introduction of the tuning parameter $\varepsilon$ guarantees $w_{k}\geq\varepsilon$ and consequently ensures that various importance ratio based estimators used in our algorithm are always well-defined.
The effective policy at round $t$ is therefore the convex mixture
$Q_{t}(a\mid x_{t})\;=\;\sum_{k=1}^{K}w_{t,k}\,\pi_{k}(a\mid x_{t}).$ (11)
The learner then draws an action $a_{t}\sim Q_{t}(\cdot\mid x_{t})$ and observes the corresponding stochastic loss $\ell_{t}$. Throughout, we work under the following assumptions:
Assumption A.
  1. (A1)
    $\mathbb{E}[\varepsilon_{t}|\mathcal{F}_{t-1},x_{t}]=0$, $|\varepsilon_{t}|\leq 1$, and $\mathbb{E}[\varepsilon^{2}_{t}|\mathcal{F}_{t-1},x_{t},a_{t}]=\sigma^{2}$ for all $t\geq 1$.
  2. (A2)
    The feature vector and the unknown parameter satisfy
    $\|\beta^{\star}\|_{2}\leq 1\quad\text{and}\quad\|c(x,a)\|_{2}\leq 1,\qquad\text{for all }(x,a)\in\mathcal{X}\times\mathcal{A}.$
  3. (A3)
    The context vectors $(x_{t})_{t\geq 1}$ are drawn i.i.d. from a distribution $\mathcal{P_{X}}$, and for every expert $k\in[K]$
    $\lambda_{\min}\left\{\mathbb{E}_{x\sim\mathcal{P_{X}}}\left(\sum_{a\in\mathcal{A}}\pi_{k}(a\mid x)\cdot c(x,a)c(x,a)^{\top}\right)\right\}\geq\lambda^{\star}>0.$ (12)
  4. (A4)
    Assume that $\sup_{k\in[K]}\{\inf_{a,x}\pi_{k}(a|x)\}>0$.
Assumption (A1) states that the noise sequence is conditionally zero mean, bounded and homoscedastic. Assumption (A2) imposes a boundedness assumption on the parameter vector $\beta^{\star}$ and feature map $c(\cdot,\cdot)$. Assumptions (A1) and (A2) together ensure that the observed loss $\{\ell_{t}\}$ is uniformly bounded. Assumption (A3) posots that the the context vectors $\{x_{t}\}$ are i.i.d., and that the weighted covariance matrix associated with each expert is non-singular. Assumption (A4) can always be satisfied by taking one of the expert is as a uniform expert $\pi_{\mathrm{unif}}(a\mid x)=1/|\mathcal{A}|$ for all action $a\in\mathcal{A},\;\;\text{and context}\;\;x\in\mathcal{X}$. This assumption along with our choice of weights $w_{tk}\geq\epsilon>$ (see equation (10)) ensures that $Q_{t}(a_{t}\mid x_{t})>0$ for all $t\geq 1$ and various importance weight based estimators are well-defined.
2.1 Linearity of loss in $w$
Let $\Delta_{K}$ be the simplex in $K$ dimension. At each round $t$, the learner maintains mixture weights $w_{t}\in\Delta_{K}$ over a fixed set of base policies $\{\pi_{k}\}_{k=1}^{K}$. The executed policy is the convex combination
$Q_{t}(a\mid x_{t})\;=\;\sum_{k=1}^{K}w_{t,k}\,\pi_{k}(a\mid x_{t}).$ (13)
This mixture induces both the sampling distribution of actions and the expected loss of the algorithm. The loss incurred at time $t$ is a function of both context $x_{t}$ and action $a_{t}$, due to which in the following discussion we rewrite $l_{t}$ as $l_{t}(a_{t},x_{t})$. Now, at round $t$, any expectation under $Q_{t}$ is a convex combination of the expectations under the individual experts $\{\pi_{k}\}$:
[…]
$\equiv\left\langle w_{t},g^{\star}(x_{t})\right\rangle$ (15)
Hence, the expected loss is linear in $w_{t}$. Since the per-round loss $g^{\star}(x_{t})$ depends on the context, we define the global (context-averaged) loss vector
$\bar{g}^{\star}:=\mathbb{E}_{x\sim\mathcal{P_{X}}}[\,g^{\star}(x)\,]=\Big(\,\mathbb{E}_{x}\!\big[\sum_{a}l(a,x)\,\pi_{k}(a\mid x)\big]\Big)_{k=1}^{K}.$ (16)
2.2 Unbiased estimate of $g^{\star}(x_{t})$
Although only one action $a_{t}$ is observed, the importance-weighted estimator for expert $k$,
$\widehat{g}_{t,k}\;=\;\,\ell_{t}\,\frac{\pi_{k}(a_{t}\mid x_{t})}{Q_{t}(a_{t}\mid x_{t})},$ (17)
is an unbiased estimator of the per-expert loss $g_{k}^{\star}(x_{t})=\sum_{a}l(a,x_{t})\pi_{k}(a\mid x_{t})$. Taking expectation over $a_{t}\sim Q_{t}(\cdot\mid x_{t})$ yields
$\mathbb{E}[\widehat{g}_{t,k}\mid x_{t}]$ $=\sum_{a}\mathbb{E}[l(a,x_{t})\mid x_{t}]\,Q_{t}(a\mid x_{t})\,\frac{\pi_{k}(a\mid x_{t})}{Q_{t}(a\mid x_{t})}=g_{k}^{\star}(x_{t}).$ (18)
2.3 Regret
We measure the regret of our policy with respect to optimal mixture:
$w^{\star}:=\arg\min_{w\in\Delta_{K}}\left\langle\bar{g}^{\star},w\right\rangle$
The cumulative regret after $T$ rounds is defined as
$\text{Reg}(T)\;=\;\sum_{t=1}^{T}\mathbb{E}\left\{\left\langle g^{\star}(x_{t}),w_{t}-w^{\star}\right\rangle\right\}.$ (19)
All expectations are taken with respect to the learner’s internal randomness and the stochastic loss environment as well as the random context $x_{t}$.
3 A Penalized Exp4 Algorithm
In this section we describe our algorithm. Recall that at round $t$, the effective policy executed by the learner is the convex mixture
$Q_{t}(a\mid x_{t})\;=\;\sum_{k=1}^{K}w_{t,k}\,\pi_{k}(a\mid x_{t}),$ (20)
where each $\pi_{k}(a\mid x_{t})$ is an arbitrary base policy that maps the context $x_{t}$ to a distribution over actions. Upon taking an action $a_{t}\sim Q_{t}(\cdot\mid x_{t})$, the learner receives a stochastic loss $\ell_{t}$. To evaluate the experts, our algorithm — stated in Algorithm 3 — constructs the standard importance-weighted gradient estimator
$\widehat{g}_{t,k}=\;\,\ell_{t}\,\frac{\pi_{k}(a_{t}\mid x_{t})}{Q_{t}(a_{t}\mid x_{t})},\qquad k=1,\ldots,K.$ (21)
The weight vector $w_{t}$ is then updated via a composite mirror-descent duchi2011adaptive step with the entropy-induced mirror map $\phi(\cdot)$ , and an entropy-induced penalty $R(\cdot)$:
$\phi(w)=\sum_{k=1}^{K}w_{k}\log w_{k}-w_{k}\qquad\text{and}\qquad R(w)=\sum_{k=1}^{K}w_{k}\left(\log w_{k}+\log(1/\epsilon)-1\right)$ (22)
Formally, our Algorithm 1 minimizes the following loss in an online fashion via a mirror descent-style algorithm:
$\min_{w\in\Delta_{\varepsilon}}\left\{\left\langle\bar{g}^{\star},w\right\rangle+\lambda R(w)\right\}.$ (23)
The regularization term $\lambda R(w)$ introduces a mild curvature to the otherwise linear objective $\langle\bar{g}^{\star},w\rangle$, ensuring better-behaved updates. The parameter $\lambda$ is chosen to be small so that the solution to the modified problem (23) achieves low regret with respect to the natural linear loss $\langle\bar{g}^{\star},w\rangle$. In addition, the constraint set $\Delta_{\varepsilon}$ guarantees that the gradients of the loss (23) remain well-behaved.
Algorithm 1 Penalized EXP4
1:Input: Base policies $\{\pi_{k}\}_{k=1}^{K}$, stepsizes $\{\eta\}$, penalty $\lambda>0$, floor $\varepsilon\in(0,1/K]$.
2: Initialize weights $w_{1,k}=1/K$ for all $k$, and set $c_{R}=\log(1/\epsilon)$
3:for $t=1,\ldots,T$ do
4:  Observe context $x_{t}$
5:  Form mixture $Q_{t}(a\mid x_{t})=\sum_{k=1}^{K}w_{t,k}\,\pi_{k}(a\mid x_{t})$
6:  Sample $a_{t}\sim Q_{t}(\cdot\mid x_{t})$ and observe loss $\ell_{t}$
7:  Compute estimates $\widehat{g}_{t,k}=\,\ell_{t}\,\pi_{k}(a_{t}\mid x_{t})/Q_{t}(a_{t}\mid x_{t})$
8:  Update intermediate weights
$w^{+}_{t+1,k}=w_{t,k}\,\exp\!\Big(-\eta\,\widehat{g}_{t,k}-\lambda\big([\nabla R(w_{t})]_{k}+c_{R}\big)\Big)$
9:  Update $w_{t+1}=\arg\min\limits_{w\in\Delta_{\varepsilon}}\left\{D_{\phi}(w,w_{t+1}^{+})\right\}$
10:end for
4 Main results
In this section we state our main results. Section 1.1 highlights that, for general adaptive data-collection rules, valid confidence intervals for linear functionals must inflate by a factor on the order of $\sqrt{d\log T}$. In this section, we show that the penalized EXP4 procedure (Algorithm 1) avoids this worst-case behavior. In particular, this algorithm satisfies the stability from Definition 1, which ensures the validity of Wald-type inference, and it does so while retaining near–minimax-optimal regret.
We begin with the stability and asymptotic normality of the ordinary least-squares estimator, and then establish regret guarantees.
4.1 Stability and a central limit theorem
A key component of our analysis is that the empirical average of the weight vectors $\{w_{t}\}$ converges to a fixed vector $w_{T}^{\star}$:
$\frac{1}{T}\sum_{t=1}^{T}w_{t,k}-w^{\star}_{T,k}\;\xrightarrow[]{P}\;0,\qquad k\in[K].$ (24)
A consequence of this convergence is that the Gram matrix $S_{T}:=\sum_{t=1}^{T}c(x_{t},a_{t})c(x_{t},a_{t})^{\top}$ is asymptotically close to a deterministic limit $\Sigma^{\star}_{T}$, where
$\Sigma^{\star}_{T}=T\sum^{K}_{k=1}w^{\star}_{T,k}\Sigma_{k}\qquad\text{with}\qquad\Sigma_{k}:=\mathbb{E}_{x\sim\mathcal{P_{X}}}\left(\sum_{a\in\mathcal{A}}\pi_{k}(a\mid x)\cdot c(x,a)c(x,a)^{\top}\right)$ (25)
Thus the stability condition (1) is satisfied. With this structural ingredient in place, we now state our quantitative central limit theorem.
Theorem 1.
Under Assumptions (A1)-(A4), the penalized-EXP4 algorithm (1) with step size $\eta=\sqrt{\frac{\log K}{|\mathcal{A}|T}}$, and tuning parameters $\epsilon=\frac{1}{KT}$, $\lambda=\frac{\gamma_{T}}{\sqrt{T}}$ with $\gamma_{T}\rightarrow\infty$ and $T/\log^{2}T\gg\gamma_{T}$, is stable; we have
$\Sigma_{T}^{\star-1}\cdot\mathrm{S}_{T}\xrightarrow{p}\mathbf{I}_{d},$
with $\Sigma_{T}^{\star}$ from (25). Consequently, for any fixed $a\in^{d}$ we have
$\frac{a^{\top}(\widehat{\beta}_{\mathrm{OLS}}-\beta^{\star})}{\widehat{\sigma}\sqrt{a^{\top}S_{T}^{-1}a}}\xrightarrow{d}\mathcal{N}\left(0,1\right)$ (26a)
where $\widehat{\beta}_{\mathrm{OLS}}$ is the least square estimate, and $\widehat{\sigma}$ is any consistent estimator of noise standard deviation $\sigma$.
Throughout we assume that the sample covariance matrix $S_{T}$ is invertible. If $S_{T}$ is not invertible, one can construct a ridge estimator with the same guarantees. We highlight the technical details in Section 6. A direct consequence of Theorem 1 is that the Wald confidence interval (4) is asymptotically exact: for any fixed $a\in\mathbb{R}^{d}$,
$\lim_{T\to\infty}\mathbb{P}\!\left(a^{\top}\beta^{\star}\in\mathcal{I}^{\mathrm{Wald}}_{T}(a)\right)\;=\;1-\alpha.$
Thus, stability ensures that the classical (nonadaptive) form of the Wald interval remains valid despite the adaptively collected data. See Section 6.1 for a proof of Theorem 1.
4.2 Regret guarantees
We next turn to regret guarantees. The following theorem quantifies how the choice of $\gamma_{T}$ influences the regret, formalizing the tradeoff introduced by the stabilizing penalty.
Theorem 2.
Suppose Assumptions (A1)-(A4) are in force. Then, for $T\geq 4$, under the stepsize $\eta=\sqrt{\frac{\log K}{TK}}$ and tuning parameters $\epsilon=\frac{1}{KT}$ and $\lambda=\frac{\gamma_{T}}{\sqrt{T}}$, Algorithm 1 satisfies
$\mathrm{Regret}(T)\;\leq\;8\sqrt{TK\log K}\;+\;\gamma_{T}\log(KT)\sqrt{T}\;+\;\frac{4\gamma_{T}^{2}\log^{3}(KT)}{K^{2}\sqrt{T}}.$ (27)
We prove Theorem 2 in Section 6.2. Setting $\gamma_{T}=\sqrt{\log T}$ yields regret of order $O\!\left(\sqrt{TK\log(KT)}\right)$, matching minimax lower bounds up to logarithmic factors. Thus, despite being explicitly regularized for inferential stability, the procedure retains the characteristic efficiency of the EXP4 family.
5 Experiments
We evaluate the empirical behavior of the proposed penalized–EXP4 algorithm in a stochastic contextual bandit environment with linear loss structure. In this section we provide the simulation details for the standard OLS estimator. Experimental details for the ridge estimator are presented in Appendix B.
5.1 Simulation Environment
The loss model follows a block–sparse linear structure
$\ell_{t}=\langle x_{t},\theta_{a_{t}}\rangle+\varepsilon_{t},$
where $\varepsilon_{t}\sim\text{Unif}(-0.1,0.1)$ and $x_{t}\in\mathbb{R}^{d_{x}}$ is a normalized Gaussian context vector with $\|x_{t}\|_{2}\leq 1$. Each arm $a\in\{1,\dots,A\}$ possesses an unknown parameter $\theta_{a}\in\mathbb{R}^{d_{x}}$, and the global coefficient vector
$\beta^{\star}=(\theta_{1},\dots,\theta_{A})\in\mathbb{R}^{Ad_{x}}$
is normalized to satisfy $\|\beta^{\star}\|_{2}\leq 1$. The learner never observes $\theta_{a}$; instead, only bandit feedback $\ell_{t}$ is revealed.
We encode actions through a sparse block feature map
$c(x,a)=\big(0,\dots,x,\dots,0\big)\in\mathbb{R}^{Ad_{x}},$
Our experiments are conducted in two settings. The first assumes experts based on softmax policies, and the second replaces them with six layer neural network policies. We outline the specifics of each configuration in the following sections.
Simulation setting with softmax experts
Each expert network produces action probabilities through a softmax map:
$\pi_{k}(a\mid x)=\frac{e^{\left\langle u_{a,k},x\right\rangle}}{\sum_{a\in\mathcal{A}}e^{\left\langle u_{a,k},x\right\rangle}}$
where, $k\in[K]$ and the entries of the weight vector $u_{a,k}$ are i.i.d. draws from $\mathcal{N}(0,12)$ distribution. In our experiments we set $A=K=5$ and $d_{x}=10$.
Simulation setting with neural experts
Unlike the previous setting, the experts now form neural policies with a six-layer ReLU architecture. The expert policy is a six-layer neural network given by
$x\;\longrightarrow\;h_{1}\;\longrightarrow\;h_{2}\;\longrightarrow\;h_{3}\;\longrightarrow\;h_{4}\;\longrightarrow\;h_{5}\;\longrightarrow\;h_{6}\;\longrightarrow\;\operatorname{softmax}(\text{logits}),$
where the hidden layers satisfy
$h_{i}=\operatorname{ReLU}\!\left(W_{i}h_{i-1}+b_{i}\right),\qquad i=1,\ldots,6,$
with $h_{0}=x$. The entries of the weight matrices $W_{i}$ are i.i.d. draws from a standard Gaussian random variable. The resulting expert policy is
$\pi(a\mid x)=\frac{\exp(\text{logits}_{a})}{\sum_{a^{\prime}=1}^{A}\exp(\text{logits}_{a^{\prime}})}.$
In our experiments we set $A=3$ and $K=5$ and $d_{x}=50$.
5.2 Algorithmic Configuration
Let $z_{t}=c(x_{t},a_{t})$. At the end of horizon $T$, we compute the OLS estimator
$\widehat{\beta}_{T}=S_{T}^{-1}b_{T},\qquad S_{T}=\sum_{t=1}^{T}z_{t}z_{t}^{\top},\quad b_{T}=\sum_{t=1}^{T}z_{t}\ell_{t}.$
Hyperparameters for Algorithm 3 are set to be
$\varepsilon=\frac{1}{KT},\qquad\lambda_{\rm pen}=\frac{\sqrt{\log T}}{\sqrt{T}},\qquad\eta=\sqrt{\frac{\log K}{|\mathcal{A}|T}},$
unless stated otherwise. We draw a random unit direction $a\in\mathbb{R}^{Ad_{x}}$ and for each confidence level $\alpha\in[0.20,0.01]$, check whether the true parameter lies inside the interval. Concretely we check if the target parameter $a^{\top}\beta^{\star}$ lies in the interval
$\mathcal{I}^{\mathrm{APS}}_{T}(a):=\big[a^{\top}\widehat{\beta}_{\mathrm{ridge}}-R_{T}\sqrt{a^{\top}\mathbf{V}_{T}^{-1}a},\;a^{\top}\widehat{\beta}_{\mathrm{ridge}}+R_{T}\sqrt{a^{\top}\mathbf{V}_{T}^{-1}a}\big],$
where $R_{T}$ is from (3) with $S=1$. Wald coverage is measured analogously using the confidence interval $\mathcal{I}^{\mathrm{Wald}}_{T}(a)$ from (4) , with $\widehat{\sigma}$ as the sample standard deviation estimate (lai1982least, Lemma 3). For each $T\in\{500,3000\}$, we report empirical coverage of $\mathcal{I}^{\mathrm{APS}}(a)$ vs $\mathcal{I}^{\mathrm{Wald}}(a)$, and their average width.
5.3 Simulation Plots
To quantify inferential quality we compute, over $N_{\mathrm{runs}}=1200$ Monte Carlo trials, the empirical coverage and interval width of confidence intervals produced by our method and a Wald-type baseline.
Softmax Experts
We vary the horizon and confidence level:
$T\in\{500,3000\},\qquad A=K=5,\qquad d_{x}=10,$
Figure 1: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Simulations are based on $T=500$ runs.
Figure 2: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Simulations are based on $T=3000$ runs.
$\frac{a^{\top}\left(\widehat{\beta}_{\mathrm{OLS}}-\beta^{\star}\right)}{\sqrt{a^{\top}S_{T}^{-1}a}}$
$\frac{a^{\top}\left(\widehat{\beta}_{\mathrm{OLS}}-\beta^{\star}\right)}{\sqrt{a^{\top}S_{T}^{-1}a}}$
Figure 3: Left: Histogram of the standardized OLS estimator for the Softmax Experts with $T=500$. Right: Histogram of the standardized OLS estimator for the Softmax Experts with $T=3000$.
Neural Experts
We vary the horizon and confidence level:
$T\in\{500,3000\},\qquad A=3,K=5,\qquad d_{x}=50,$
Figure 4: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Simulations are based on $T=500$ runs.
Figure 5: Left: Coverages of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Simulations are based on $T=3000$ runs.
$\frac{a^{\top}\left(\widehat{\beta}_{\mathrm{OLS}}-\beta^{\star}\right)}{\sqrt{a^{\top}S_{T}^{-1}a}}$
$\frac{a^{\top}\left(\widehat{\beta}_{\mathrm{OLS}}-\beta^{\star}\right)}{\sqrt{a^{\top}S_{T}^{-1}a}}$
Figure 6: Left: Histogram of the standardized OLS estimator for the 66-layer Neural-Network Experts with $T=500$. Right: Histogram of the standardized OLS estimator for the 66-layer Neural-Network Experts with $T=3000$.
6 Proof of Theorems
In this section we prove Theorems 1 and 2. Recall that $S_{T}=\sum^{T}_{t=1}c(x_{t},a_{t})c(x_{t},a_{t})^{\top}$ where, the random vector $c(x,a)\in^{d}$ satisfies
$\sup_{x,a}\|c(x,a)\|_{2}<\infty.$ (28)
By stability (definition (1)) we mean that the sample covariance matrix satisfies the following property.
${\Sigma^{\star}_{T}}^{-1}S_{T}\;\stackrel{{\scriptstyle p}}{{\rightarrow}}\;\mathbf{I}\qquad\text{where}\;\;\{{\Sigma^{\star}_{T}}\}\;\;\text{is a sequence of non-random positive definite matrices.}$ (29)
Let us define the matrix
$Y_{t}:=c(x_{t},a_{t})c(x_{t},a_{t})^{\top}.$
We denote by $\mathcal{F}_{t}:=\sigma(x_{1},a_{1},r_{1},\ldots,x_{t},a_{t},\ell_{t})$, the $\sigma$-field generated by all observations up to time $t$, including $x_{t}$, $a_{t}$ and $\ell_{t}$. Since $x_{t}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{P_{X}}$, the conditional expectation
$\mathbb{E}[Y_{t}\mid\mathcal{F}_{t-1}]=\mathbb{E}_{x_{t}}\!\left[\sum_{a\in\mathcal{A}}Q_{t}(a\mid x_{t})\,c(x_{t},a)c(x_{t},a)^{\top}\right]=\sum^{K}_{k=1}w_{t,k}\Sigma_{k}$
where the matrices
$\Sigma_{k}:=\mathbb{E}_{x}\!\left[\sum_{a\in\mathcal{A}}\pi_{k}(a\mid x)\,c(x,a)c(x,a)^{\top}\right],\qquad k=1,\ldots,K$ (30)
are population-level second moments under each expert. In order to simplify notations, throughout we assume $\sigma=1$.
6.1 Proof of Theorem 1
We first prove stability holds for the standard OLS estimator, assuming that $S_{T}$ is invertible, followed by the proof for the modified ridge estimator. Once the stability condition (1) is verified, the asymptotic normality claim (26a) follows directly by invoking Theorem 3 from lai1982least along with an application of Slutsky’s theorem.
Part (a): proof of stability condition for $S_{T}$ (29)
In order to prove the stability of the sample covariance matrix $S_{T}$, we need to first prove that the average of the random weight vectors $\{w_{t}\}_{t\geq 0}$ converge to a non-random vector, which we call $w_{T}^{\star}$. Concretely,
$\frac{1}{T}\sum_{t=1}^{T}w_{t,k}-w^{\star}_{T,k}\;\stackrel{{\scriptstyle p}}{{\rightarrow}}\;0$ (31)
for all experts $k\in[K]$.
Proof of weight stability (31):
We begin by recalling that $\widetilde{g}_{t}=\widehat{g}_{t}+\lambda\nabla R(w_{t})$. Let $\tilde{R}_{\lambda}(w):=\lambda R(w)+\langle\bar{g}^{\star},w\rangle$. If $f$ is any $M$-strongly convex function, then $f+g$ is also $M$-strongly convex if $g$ is a linear function (nesterov2013introductory). This result implies that $\tilde{R}_{\lambda}$ is $\lambda$ strongly convex. Now, for any arbitrary $y\in\Delta_{\epsilon}$ we have,
$\dfrac{1}{T}\sum_{t=1}^{T}\mathbb{E}\langle\tilde{g}_{t},w_{t}-y\rangle$ $=\dfrac{1}{T}\sum_{t=1}^{T}\mathbb{E}\langle\bar{g}^{\star}+\lambda\nabla R(w_{t}),w_{t}-y\rangle$
$=\dfrac{1}{T}\sum_{t=1}^{T}\bigg(\mathbb{E}\langle\bar{g}^{\star},w_{t}-y\rangle+\lambda\mathbb{E}\langle\nabla R(w_{t}),w_{t}-y\rangle\bigg)$
$\overset{(i)}{\geq}\dfrac{1}{T}\sum_{t=1}^{T}\bigg(\mathbb{E}\langle\bar{g}^{\star},w_{t}-y\rangle+\lambda\mathbb{E}(R(w_{t})-R(y))\bigg)$
$\geq\dfrac{1}{T}\sum_{t=1}^{T}\bigg(\mathbb{E}[\langle\bar{g}^{\star},w_{t}\rangle+\lambda R(w_{t})]-\mathbb{E}[\langle\mu,y\rangle+\lambda R(y)]\bigg)$
$\overset{(ii)}{=}\dfrac{1}{T}\sum_{t=1}^{T}\mathbb{E}\bigg[\tilde{R}(w_{t})-\tilde{R}(y)\bigg]$
$\overset{(iii)}{\geq}\mathbb{E}\bigg[\tilde{R}(\bar{w}_{T})-\tilde{R}(y)\bigg]$
$\overset{(iv)}{=}\mathbb{E}\bigg[D_{\tilde{R}_{\lambda}}(\bar{w}_{T},y)+\langle\nabla\tilde{R}_{\lambda}(y),\bar{w}_{T}-y\rangle)\bigg]$ (32)
Where $D_{\tilde{R}_{\lambda}}$ is the Bregman divergence defined in terms of $\tilde{R}_{\lambda}$. The first equality above follows by the unbiasedness of the gradient estimate $\widehat{g}_{t}$ (Section 2.2). Equation $(i)$ above follows from convexity of $R$, equation $(ii)$ follows from definition of $\tilde{R}$ and $(iii)$ holds due to convexity of $\tilde{R}_{\lambda}$. Equality $(iv)$ follows from the definition of Bregman divergence. Note that the above discussion holds for any $y\in\Delta_{\epsilon}$. Let us define $w^{\star}_{T}$ as follows
$w^{\star}_{T}=\arg\min_{w\in\Delta_{\epsilon}}\tilde{R}_{\lambda}(w)$ (33)
If we choose $y=w^{\star}_{T}$ then by applying the first order optimality criteria of equation (33) and Pinsker’s inequality we have
$\langle\nabla\tilde{R}_{\lambda}(w^{\star}_{T}),\bar{w}_{T}-w^{\star}_{T}\rangle\geq 0\quad\text{and,}\ \ D_{\tilde{R}_{\lambda}}(\bar{w}_{T},w^{\star}_{T})\geq\dfrac{\lambda}{2}\|\bar{w}_{T}-w^{\star}_{T}\|_{1}^{2}$ (34)
From equations (6) and (34), we arrive at the following lower bound:
$\dfrac{1}{T}\sum_{t=1}^{T}\mathbb{E}\langle\tilde{g}_{t},w_{t}-w^{\star}_{T}\rangle\geq\dfrac{\lambda}{2}\mathbb{E}\|\bar{w}_{T}-w^{\star}_{T}\|_{1}^{2}$ (35)
Now, the following lemma provides an upper bound on $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\langle\tilde{g}_{t},w_{t}-w^{\star}_{T}\rangle$.
Lemma 1.
Under the setup of Theorem 1, we have
$\dfrac{1}{T}\sum_{t=1}^{T}\mathbb{E}\langle\tilde{g}_{t},w_{t}-w^{\star}_{T}\rangle\leq 5{\frac{\sqrt{|\mathcal{A}|\log K}}{\sqrt{T}}}+\frac{4\gamma_{T}^{2}\log(K)\log^{2}\left(KT\right)}{|\mathcal{A}|^{2}\sqrt{T}}$
We prove this lemma in Section 6.2 . From equation (35) and Lemma 1 it follows from Jensen’s inequality that,
$\mathbb{E}[\|\bar{w}_{T}-w^{\star}_{T}\|_{1}]$ $\leq\sqrt{\frac{2}{\lambda_{T}}}\sqrt{\dfrac{1}{T}\sum_{t=1}^{T}\mathbb{E}\langle\tilde{g}_{t},w_{t}-w^{\star}_{T}\rangle}$
$\leq\sqrt{10{\frac{\sqrt{|\mathcal{A}|\log K}}{\sqrt{T}\lambda_{T}}}+\frac{8\gamma_{T}^{2}\log(K)}{\lambda_{T}|\mathcal{A}|^{2}T\sqrt{T}}\log^{2}\left(KT\right)}$
$=\sqrt{10{\frac{\sqrt{|\mathcal{A}|\log K}}{\gamma_{T}}}+\frac{8\gamma_{T}\log(K)\log^{2}\left(KT\right)}{|\mathcal{A}|^{2}T}}$
Hence, if $|\mathcal{A}|$ and $K$ do not vary with $T$ then for any $\gamma_{T}\rightarrow\infty$ such that $T/\log^{2}T\gg\gamma_{T}$ it follows that $\frac{1}{T}\sum_{t=1}^{T}w_{t,k}-w^{\star}_{T,k}\;\stackrel{{\scriptstyle p}}{{\rightarrow}}\;0$ for all experts $k\in[K]$.
Now we are ready to prove stability of $S_{T}$. We first decompose $S_{T}$ into two parts:
$S_{T}=M_{T}+B_{T}\qquad\text{where}$ (36a)
$D_{t}:=Y_{t}-\mathbb{E}[Y_{t}\mid\mathcal{F}_{t-1}],\;\;M_{T}:=\sum_{t=1}^{T}D_{t},\;\;\text{and}\;\;B_{T}:=\sum_{t=1}^{T}\mathbb{E}[Y_{t}\mid\mathcal{F}_{t-1}].$ (36b)
Now, we define
$\frac{1}{T}{\Sigma^{\star}_{T}}:=\sum^{K}_{k=1}w^{\star}_{T,k}\Sigma_{k}$ (37)
where
$\Sigma_{k}:=\mathbb{E}_{x\sim\mathcal{P_{X}}}\left(\sum_{a\in\mathcal{A}}\pi_{k}(a\mid x)\cdot c(x,a)c(x,a)^{\top}\right)$
We show, using Lemma 2 stated below, that
$\dfrac{S_{T}}{T}=\dfrac{M_{T}}{T}+\dfrac{B_{T}}{T}=\underbrace{\dfrac{M_{T}}{T}}_{o_{\mathbb{P}}(1)}+\underbrace{\left[\dfrac{B_{T}}{T}-\frac{1}{T}{\Sigma^{\star}_{T}}\right]}_{o_{\mathbb{P}}(1)}+\frac{1}{T}{\Sigma^{\star}_{T}}$ (38)
Lemma 2.
Let $\Sigma_{k},M_{T}$ and $B_{T}$ be as defined in equations (36a). Suppose that there exists non-random weight vector $w^{\star}_{T}=(w^{\star}_{T,1},\ldots,w^{\star}_{T,K})^{\top}$ such that $\frac{1}{T}\sum_{t=1}^{T}w_{t,k}-w^{\star}_{T,k}\;\stackrel{{\scriptstyle p}}{{\rightarrow}}\;0$ for all experts $k\in[K]$. Then we have,
[…]
We prove this Lemma in Appendix C. Let $\lambda^{\star}$ be the minimum eigenvalue among $\lambda_{\min}(\Sigma_{k})$ for all $k\in[K]$. Note that $\lambda_{\min}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)\geq\lambda^{\star}>0$ by Assumption (A3). Hence, the minimum eigenvalue of $\Sigma_{T}^{\star}/T$ — which is a convex combination of $\Sigma_{k}^{\prime}s$ — is also lower bounded by $\lambda^{\star}$. This fact combined with the decomposition (38) yields
${\Sigma^{\star}_{T}}^{-1}S_{T}\xrightarrow{\mathbb{P}}\mathbf{I}.$
Part (b) : Proof of stability condition for ridge estimator
In this section we show that if we consider a ridge estimator, with penalty term $\lambda_{rid}\ll\sqrt{T}$, then Theorem 1 holds. Let $R_{T}:=S_{T}+\lambda_{rid}\ \mathbf{I}$ and consider the ridge estimator $\hat{\beta}_{rid}:=R_{T}^{-1}S_{T}\ \beta^{\star}+\ R_{T}^{-1}\sum^{T}_{t=1}z_{t}\varepsilon_{t}$, where $z_{t}=c(x_{t},a_{t})$. Our statistic of interest is the following
$\mathcal{R}_{1}(T):=\dfrac{a^{\top}(\hat{\beta}_{rid}-\beta^{\star})}{\sqrt{a^{\top}R_{T}^{-1}a}}$ (40)
We are interested to show that for the design matrix $R_{T}=S_{T}+\lambda_{rid}\mathbf{I}$, there exists a sequence of deterministic, positive definite matrices $\{\tilde{\Sigma}_{T}\}$ such that
$\tilde{\Sigma}^{-1}_{T}R_{T}\xrightarrow{\mathbb{P}}\mathbf{I}$ (41)
Now, if we choose our penalty term $\lambda_{rid}\ll T$ then for the ridge estimator we have the decomposition of $R_{T}$ :
$\dfrac{R_{T}}{T}=\underbrace{\dfrac{M_{T}}{T}}_{o_{\mathbb{P}}(1)}+\underbrace{\left[\dfrac{B_{T}}{T}-\frac{1}{T}{\Sigma^{\star}_{T}}\right]}_{o_{\mathbb{P}}(1)}+\underbrace{\frac{\lambda_{rid}}{T}}_{o_{\mathbb{P}}(1)}\mathbf{I}+\frac{1}{T}{\Sigma^{\star}_{T}}$ (42)
Therefore, applying Lemma 2 along with the fact that $\lambda_{\min}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)>0$ shows that property (41) is satisfied by choosing $\tilde{\Sigma}_{T}={\Sigma^{\star}_{T}}$, where $\Sigma^{\star}_{T}$} is as defined in equation (37).
By substituting the value of $\hat{\beta}_{rid}$ in equation (40), we obtain the following decomposition
$\mathcal{R}_{1}(T)=\underbrace{\dfrac{a^{\top}(R_{T}^{-1}S_{T}-\mathbf{I})}{\sqrt{a^{\top}R_{T}^{-1}a}}\beta^{\star}}_{\mathcal{T}_{1}(T)}\ \ +\ \ \underbrace{\dfrac{a^{\top}R_{T}^{-1}}{\sqrt{a^{\top}R_{T}^{-1}a}}\sum^{T}_{t=1}z_{t}\varepsilon_{t}}_{\mathcal{T}_{2}(T)}$ (43)
The term $\mathcal{T}_{1}(T)$ in equation (43) is the bias induced by the regularization. As $R_{T}:=S_{T}+\lambda_{rid}\ \mathbf{I}$ we note that $R_{T}^{-1}S_{T}-\mathbf{I}=R_{T}^{-1}(S_{T}-R_{T})$, which is equal to $\lambda_{rid}\ R_{T}^{-1}$. Therefore, $\mathcal{T}_{1}(T)$ can be rewritten as
$\mathcal{T}_{1}(T)=-\lambda_{rid}\dfrac{a^{\top}R_{T}^{-1}}{\sqrt{a^{\top}R_{T}^{-1}a}}\beta^{\star}$
Let $k_{T}:=\sqrt{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}a}/\sqrt{a^{\top}R_{T}^{-1}a}$ and $\mathcal{M}_{T}=k_{T}\times R_{T}^{-1}{\Sigma^{\star}_{T}}$. Then,
$\mathcal{T}_{1}(T)=\frac{-\lambda_{rid}}{\sqrt{T}}\dfrac{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}}{\sqrt{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}a}}\beta^{\star}+\frac{-\lambda_{rid}}{\sqrt{T}}\dfrac{a^{\top}[\mathcal{M}_{T}-\mathbf{I}]\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}}{\sqrt{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}a}}\beta^{\star}$
Therefore, as $\lambda_{\min}(\frac{1}{T}{\Sigma^{\star}_{T}})$ is uniformly bounded away from zero and $\lambda_{\max}(\frac{1}{T}{\Sigma^{\star}_{T}})$ is uniformly bounded above, it follows from stability (equation (41)) that by choosing $\lambda_{rid}\ll\sqrt{T}$, $\mathcal{M}_{T}\xrightarrow{\mathbb{P}}\mathbf{I}$ and consequently, the bias $\mathcal{T}_{1}(T)$ converges to 0 in probability, as $T\rightarrow\infty$.
Now, for $\mathcal{T}_{2}(T)$ we have a similar decomposition:
$\mathcal{T}_{2}(T)=\underbrace{\dfrac{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}}{\sqrt{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}a}}\frac{1}{\sqrt{T}}\sum^{T}_{t=1}z_{t}\varepsilon_{t}}_{I_{1}(T)}\ \ +\ \ \underbrace{\dfrac{a^{\top}[\mathcal{M}_{T}-\mathbf{I}]\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}}{\sqrt{a^{\top}\left(\frac{1}{T}{\Sigma^{\star}_{T}}\right)^{-1}a}}\frac{1}{\sqrt{T}}\sum^{T}_{t=1}z_{t}\varepsilon_{t}}_{I_{2}(T)}$
Now, from central limit theorem for martingale difference sequence (dvoretzky1972asymptotic), $I_{1}(T)\xrightarrow{d}\mathcal{N}(0,1)$ . This result, combined with the fact that $\mathcal{M}_{T}\xrightarrow{\mathbb{P}}\mathbf{I}$ implies $I_{2}(T)$ converges to 0 in probability. This completes the proof.
6.2 Proof of Theorem 2
We first restrict the comparator $w$ to the truncated simplex. Without loss of generality we assume arm $1$ is among the optimal arms. Now define
$w^{\star}=(1,0,\ldots,0)\qquad\text{and}\qquad w_{\epsilon}=\big(1-(K-1)\varepsilon,\;\varepsilon,\ldots,\varepsilon\big)\in\Delta_{\varepsilon}.$
To prove this theorem, we shall apply the master equation stated below.
Lemma 3.
For any sequence of contexts $\{x_{t}\}_{t=1}^{T}$ and stochastic losses $\{\ell_{t}\}_{t=1}^{T}$, the outputs $\{w_{t}\}_{t\geq 1}$ produced by Algorithm 1 satisfy, for any $y\in\Delta_{\epsilon}$
[…]
By construction, $w^{\star}\in\arg\min_{w}\langle\bar{g}^{\star},w\rangle$, and $w_{\epsilon}\in\Delta_{\epsilon}$. Consequently, we can apply Lemma 3 with the comparator $w=w_{\epsilon}$. An application of Holder’s inequality and using $\|\bar{g}^{\star}\|_{\infty}\leq 2$ now yields
$\mathrm{Regret}(T)$ $:=\sum_{t=1}^{T}\mathbb{E}\langle\bar{g}^{\star},w_{t}-w^{\star}\rangle$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\sum_{t=1}^{T}\mathbb{E}\langle\bar{g}^{\star},w_{t}-w_{\epsilon}\rangle+4TK\epsilon$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\sum_{t=1}^{T}\mathbb{E}\langle\tilde{g}_{t},w_{t}-w_{\epsilon}\rangle+4TK\epsilon+2\lambda T\log(1/\epsilon)$ (45)
Inequality $(i)$ utilizes the bound $|\langle\bar{g}^{\star},w^{\star}-w_{\epsilon}\rangle|\leq\|\bar{g}^{\star}\|_{\infty}\|w^{\star}-w_{\epsilon}\|_{1}\leq 4K\epsilon$; inequality $(ii)$ utilizes the relation
$\mathbb{E}\langle\bar{g}^{\star},w_{t}-w_{\epsilon}\rangle$ $=\mathbb{E}\langle\hat{g}_{t},w_{t}-w_{\epsilon}\rangle$
$=\mathbb{E}\langle\tilde{g}_{t},w_{t}-w_{\epsilon}\rangle-\lambda\mathbb{E}\langle\nabla R(w_{t}),w_{t}-w_{\epsilon}\rangle$
$\leq\mathbb{E}\langle\tilde{g}_{t},w_{t}-w_{\epsilon}\rangle+\lambda|\mathbb{E}\langle\nabla R(w_{t}),w_{t}-w_{\epsilon}\rangle|$
$\leq\mathbb{E}\langle\tilde{g}_{t},w_{t}-w_{\epsilon}\rangle+\lambda\|\nabla R(w_{t})\|_{\infty}\mathbb{E}[\|w_{t}-w_{\epsilon}\|_{1}]$
and the bound $\|\nabla R(w)\|_{\infty}\leq\log(1/\epsilon)$ for all $w\in\Delta_{\epsilon}$ along with Holder’s inequality. It now remains to bound the inner product term $\sum_{t=1}^{T}\mathbb{E}\langle\tilde{g}_{t},w_{t}-w_{\epsilon}\rangle$ using Lemma 3. This result is provided by Lemma 1 in the proof of the stability property of $S_{T}$. For the sake of completeness, we state the lemma below.
Lemma.
Under the setup of Theorem 1, we have
$\dfrac{1}{T}\sum_{t=1}^{T}\mathbb{E}\langle\tilde{g}_{t},w_{t}-w^{\star}_{T}\rangle\leq 5{\frac{\sqrt{|\mathcal{A}|\log K}}{\sqrt{T}}}+\frac{4\gamma_{T}^{2}\log(K)\log^{2}\left(KT\right)}{|\mathcal{A}|^{2}\sqrt{T}}$
The proof of Lemma 1 is as follows. Invoking Lemma 3 with $y=w_{\epsilon}$ and setting $\eta=\sqrt{\frac{\log K}{|\mathcal{A}|T}}$, $\lambda=\frac{\gamma_{T}}{\sqrt{T}}$, and $w_{1}=(1/K,\ldots,1/K)$ we have
[…]
Inequality $(i)$ above utilizes $D(w,w_{1})\leq\log K$ for all $w\in\Delta_{K}$. This claim holds because
$D(w,w_{1})=\sum^{K}_{k=1}w_{k}\log\left(\frac{w_{k}}{w_{1,k}}\right)\ \leq\ \log K\sum^{K}_{k=1}w_{k}=\log K$
Inequality $(ii)$ uses that for any $w\in\Delta_{\epsilon}$
$\|\nabla R(w)\|^{2}_{w,*}:=\sum_{k=1}^{K}w_{k}\log^{2}\!\left(\frac{w_{k}}{\varepsilon}\right)\;\leq\;\sum_{k=1}^{K}w_{k}\log^{2}\!\left(\frac{1}{\varepsilon}\right)=\log^{2}\!\left(\frac{1}{\varepsilon}\right).$
[…]
Lemma 4.
Under the setup of Theorem 1 we have
[…]
The proof of Lemma 4 utilizes the property of the local norm $\|\cdot\|_{w_{t},\star}$ and the specific form of the gradient estimate $\widehat{g}_{t}$ from (17). The complete proof is provided in Appendix C. Now, by substituting equation (47) and $\varepsilon=1/KT$ in equation (46) we have
$\sum_{t=1}^{T}\mathbb{E}\langle\tilde{g}_{t},w_{t}-w_{\epsilon}\rangle\leq 5\sqrt{T|\mathcal{A}|\log K}+\frac{4\gamma^{2}_{T}\log(K)}{|\mathcal{A}|^{2}\sqrt{T}}\log^{2}\left(KT\right)$
Combining bounds (45), (46) and (47) we have
$\mathrm{Regret}(T)$ $\leq 5\sqrt{T|\mathcal{A}|\log K}+\frac{4\gamma^{2}_{T}\log(K)}{|\mathcal{A}|^{2}\sqrt{T}}\log^{2}\left(1/\epsilon\right)+4TK\epsilon+2\lambda T\log(1/\epsilon)$
Finally, substituting $\epsilon=\frac{1}{KT},\ \lambda=\frac{\gamma_{T}}{\sqrt{T}}$ and using $T\geq 4$ yields
$\mathrm{Regret}(T)\leq 8\sqrt{T|\mathcal{A}|\log K}+\gamma_{T}\log(KT)\sqrt{T}+\frac{4\gamma^{2}_{T}\log^{3}(KT)}{|\mathcal{A}|^{2}\sqrt{T}}$
7 Conclusion
We studied statistical inference in linear contextual bandits under adaptive data collection, with a focus on the validity of classical Wald-type confidence intervals. By introducing a penalized variant of the EXP4 algorithm, we demonstrated that it is possible to simultaneously enforce the Lai–Wei stability condition and achieve regret guarantees that are minimax optimal up to logarithmic factors. As a consequence, ordinary least-squares estimators satisfy a central limit theorem, and Wald confidence intervals for linear functionals are asymptotically valid without incurring the $\sqrt{d\log T}$ price of adaptivity. Together, these results establish that stability and statistical efficiency are not fundamentally incompatible in contextual bandit problems, and that careful algorithmic design can recover classical inferential guarantees even under adaptive sampling.
Several directions remain open for future work. It would be of interest to extend the present analysis to settings with growing feature dimension or fully adaptive contexts. Understanding whether analogous stability–regret tradeoffs can be achieved remains an important and challenging question.
Acknowledgments
This work was partially supported by the National Science Foundation Grant DMS-2311304 to Koulik Khamaru.
Appendix A Proof of important results
In this section we provide the proof of Lemma 3.
A.1 Proof of Lemma 3
We use the shorthand $\widetilde{g}_{t}=\widehat{g}_{t}+\lambda\nabla R(w_{t})$ as the gradient estimate of the penalized loss function at $w_{t}$. We have
$\nabla\phi(w^{+}_{t+1})=\nabla\phi(w_{t})-\eta\widetilde{g}_{t}$ (48)
We begin by analyzing the quantity $\langle\eta\widetilde{g}_{t},w_{t}-y\rangle$.
$\langle\eta\widetilde{g}_{t},w_{t}-y\rangle$ $=\langle\nabla\phi(w_{t})-\nabla\phi(w^{+}_{t+1}),w_{t}-y\rangle$
[…]
Lemma 5.
For our choice of $\phi$ the following result holds for any $x,y\in\Delta_{\epsilon}$
$D_{\phi}(x,y)=D_{\phi^{\star}}(\nabla\phi(y),\nabla\phi(x))$
where $\phi^{\star}$ is the Fenchel dual of $\phi$.
From Lemma 5 and the definition of Bregman divergence we have,
[…]
$=\phi^{\star}(\nabla\phi(w^{+}_{t+1}))-\phi^{\star}(\nabla\phi(w_{t}))-\langle\nabla\phi^{\star}(\nabla\phi(w_{t})),\nabla\phi(w^{+}_{t+1})-\nabla\phi(w_{t})\rangle$ (49)
Furthermore, the coordinate-wise positivity of $\widetilde{g}_{t}$ in (48) yields $\nabla\phi(w^{+}_{t+1})\leq\nabla\phi(w_{t})$. Hence, by applying the second order Taylor expansion in equation (A.1), we have that for an intermediate point $z_{t}=\delta\ \nabla\phi(w^{\star}_{t+1})+(1-\delta)\nabla\phi(w_{t})$,
$D_{\phi}(w_{t},w^{+}_{t+1})=\frac{1}{2}\left(\nabla\phi(w^{+}_{t+1})-\nabla\phi(w_{t})\right)^{T}\left[\nabla^{2}{\phi^{\ast}}(z_{t})\right]\left(\nabla\phi(w^{+}_{t+1})-\nabla\phi(w_{t})\right)$
where $\delta\in(0,1)$ and $\phi^{\ast}(y):=\sup_{x\in R^{d}_{>0}}\left\{\langle y,x\rangle-\phi(x)\right\}=\sum_{i=1}^{d}\exp\left\{y_{i}\right\}$ is the dual map corresponding to $\phi$. Note that $\nabla^{2}\phi^{\ast}(y)=\text{Diag}\left(\exp\{y_{1}\},\ldots,\exp\{y_{d}\}\right)$. By combining this identity with the fact that $z_{t}\leq\nabla\phi(w_{t})$ coordinate-wise we get,
$\nabla^{2}{\phi^{\ast}}(z_{t})\preceq\nabla^{2}{\phi^{\ast}}(\nabla\phi(w_{t}))=\nabla^{2}{\phi^{*}}\left(\log w_{t}\right)=\text{Diag}(w_{1},\ldots,w_{d})$
As $\|z_{t}\|_{\infty}\leq 1$ we have
[…]
Overall, we arrive at
[…]
On taking full-expectation, the left hand side is bounded below by $\eta\mathbb{E}[f(w_{t})-f(y)]$. Taking the average over $t=1$ to $T$ and using the fact that $f$ is convex yields
[…]
This completes the proof.
Appendix B Numerical Experiments for Ridge Estimator
In this section we evaluate the empirical behavior of the proposed penalized–EXP4 algorithm in a stochastic contextual bandit environment for the modified ridge estimator.
Model and Data Generation
The loss model follows a block–sparse linear structure
$\ell_{t}=\langle x_{t},\theta_{a_{t}}\rangle+\varepsilon_{t},$
where $\varepsilon_{t}\sim\text{Unif}(-0.1,0.1)$ and $x_{t}\in\mathbb{R}^{d_{x}}$ is a normalized Gaussian context vector with $\|x_{t}\|_{2}\leq 1$. Each arm $a\in\{1,\dots,A\}$ possesses an unknown parameter $\theta_{a}\in\mathbb{R}^{d_{x}}$, and the global coefficient vector
$\beta^{\star}=(\theta_{1},\dots,\theta_{A})\in\mathbb{R}^{Ad_{x}}$
is normalized to satisfy $\|\beta^{\star}\|_{2}\leq 1$. The learner never observes $\theta_{a}$; instead, only bandit feedback $\ell_{t}$ is revealed. We encode actions through a sparse block feature map
$c(x,a)=\big(0,\dots,x,\dots,0\big)\in\mathbb{R}^{Ad_{x}},$
Let $R_{T}=S_{T}+\lambda_{rid}\mathbf{I}$, where $S_{T}$ is the sample covariance matrix. Here we consider the ridge estimator
$\widehat{\beta}_{rid}:=R_{T}^{-1}\sum^{T}_{t=1}c(x_{t},a_{t})l_{t}$
We draw a random unit direction $a\in\mathbb{R}^{Ad_{x}}$ and for each confidence level $\alpha\in[0.20,0.01]$, check whether the true parameter lies inside the interval. Concretely we check if the target parameter $a^{\top}\beta^{\star}$ lies in the interval
$\mathcal{I}^{\mathrm{APS}}_{T}(a):=\big[a^{\top}\widehat{\beta}_{\mathrm{ridge}}-R_{T}(\alpha)\sqrt{a^{\top}\mathbf{V}_{T}^{-1}a},\;a^{\top}\widehat{\beta}_{\mathrm{ridge}}+R_{T}(\alpha)\sqrt{a^{\top}\mathbf{V}_{T}^{-1}a}\big],$
where $R_{T}(\alpha)$ is from (3) with $S=1$. Wald coverage is measured analogously using the confidence interval $\mathcal{I}^{\mathrm{Wald}}_{T}(a)$ similar to (4),
$\mathcal{I}^{\mathrm{Wald}}_{T}(a)\;:=\;\Bigl[a^{\top}\widehat{\beta}_{rid}\;\pm\;z_{1-\alpha/2}\,\widehat{\sigma}\,\sqrt{a^{\top}R_{T}^{-1}a}\Bigr],$ (51)
with $\widehat{\sigma}$ as the sample standard deviation estimate (lai1982least, Lemma 3). For each $T\in\{500,3000\}$, we report empirical coverage of $\mathcal{I}^{\mathrm{APS}}(a)$ vs $\mathcal{I}^{\mathrm{Wald}}(a)$, and their average width.
Algorithmic Configuration
We consider the same simulation environment with $6$ layer neural network based experts as considered in Section 5. The experts now form neural policies with a six-layer ReLU architecture. The expert policy is a six-layer neural network given by
$x\;\longrightarrow\;h_{1}\;\longrightarrow\;h_{2}\;\longrightarrow\;h_{3}\;\longrightarrow\;h_{4}\;\longrightarrow\;h_{5}\;\longrightarrow\;h_{6}\;\longrightarrow\;\operatorname{softmax}(\text{logits}),$
where the hidden layers satisfy
$h_{i}=\operatorname{ReLU}\!\left(W_{i}h_{i-1}+b_{i}\right),\qquad i=1,\ldots,6,$
with $h_{0}=x$. The entries of the weight matrices $W_{i}$ are i.i.d. draws from $\mathcal{N}(0,12)$ distribution. The resulting expert policy is
$\pi(a\mid x)=\frac{\exp(\text{logits}_{a})}{\sum_{a^{\prime}=1}^{A}\exp(\text{logits}_{a^{\prime}})}.$
In our experiments we set $A=3$ and $K=5$ and $d_{x}=50$. At the end of horizon $T$, we compute the ridge estimator
$\widehat{\beta}_{rid}=\big(S_{T}+\lambda_{rid}I\big)^{-1}b_{T},\qquad S_{T}=\sum_{t=1}^{T}c(x_{t},a_{t})c(x_{t},a_{t})^{\top},\quad b_{T}=\sum_{t=1}^{T}c(x_{t},a_{t})\ell_{t}.$
Hyperparameters are selected as
$\varepsilon=\frac{1}{KT},\qquad\lambda_{\rm pen}=\frac{\sqrt{\log T}}{\sqrt{T}},\qquad\eta=\sqrt{\frac{\log K}{|\mathcal{A}|T}},\qquad\lambda_{rid}=\frac{1}{T}$
unless stated otherwise. For each $T$, we report:
  • Empirical coverage $\mathcal{I}^{\mathrm{APS}}$ vs $\mathcal{I}^{\mathrm{Wald}}$,
  • Average width $\mathcal{I}^{\mathrm{APS}}$ vs $\mathcal{I}^{\mathrm{Wald}}$,
  • Dependence on nominal confidence level $\alpha$.
Simulation Plots
Figure 7: Left: Coverage of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Simulations are based on $T=3000$ runs.
Figure 8: Left: Coverage of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Right: Expected confidence width of both $\mathcal{I}^{\mathrm{APS}}$ and $\mathcal{I}^{\mathrm{Wald}}$. Simulations are based on $T=3000$ runs.
$\frac{a^{\top}\left(\hat{\beta}_{rid}-\beta^{\star}\right)}{\sqrt{a^{\top}S_{T}^{-1}a}}$
$\frac{a^{\top}\left(\hat{\beta}_{rid}-\beta^{\star}\right)}{\sqrt{a^{\top}S_{T}^{-1}a}}$
Figure 9: Left: Histogram of the standardized Ridge estimator for $T=500$. Right: Histogram of the standardized Ridge estimator for $T=3000$.
Appendix C Auxiliary Lemmas
In this section we present some well-known results on Bregman divergence for the sake of completeness, followed by the proofs of Lemma 2 and 4.
Some Results on Bregman Divergence
We begin with two standard identities for Bregman divergences that will be used repeatedly. Let $\phi:\mathcal{D}\to\mathbb{R}$ be a differentiable, strictly convex function on an open convex set $\mathcal{D}\subset\mathbb{R}^{K}$. The Bregman divergence generated by $\phi$ is
$D_{\phi}(x,y)\;=\;\phi(x)-\phi(y)-\langle\nabla\phi(y),\,x-y\rangle,\qquad x,y\in\mathcal{D}.$
Lemma 6 (Bregman three-point identity).
For any differentiable and strictly convex function $\phi:\mathcal{D}\to\mathbb{R}$ and any $x,x^{+},y\in\mathcal{D}$,
$\big\langle\nabla\phi(x)-\nabla\phi(x^{+}),\,x-y\big\rangle\;=\;D_{\phi}(y,x)\;-\;D_{\phi}(y,x^{+})\;+\;D_{\phi}(x,x^{+}),$
where $D_{\phi}(u,v)=\phi(u)-\phi(v)-\langle\nabla\phi(v),\,u-v\rangle$ is the Bregman divergence generated by $\phi$.
Proof.
By the definition of the Bregman divergence,
$D_{\phi}(y,x)-D_{\phi}(y,x^{+})+D_{\phi}(x,x^{+})$
$=\big(\phi(y)-\phi(x)-\langle\nabla\phi(x),y-x\rangle\big)-\big(\phi(y)-\phi(x^{+})-\langle\nabla\phi(x^{+}),y-x^{+}\rangle\big)$
$\quad\;\;+\big(\phi(x)-\phi(x^{+})-\langle\nabla\phi(x^{+}),x-x^{+}\rangle\big).$
The $\phi(\cdot)$ terms cancel, giving
$\langle\nabla\phi(x),y-x\rangle+\langle\nabla\phi(x^{+}),y-x^{+}\rangle-\langle\nabla\phi(x^{+}),x-x^{+}\rangle.$
Since $y-x^{+}=(y-x)+(x-x^{+})$, the last two inner products combine to $\langle\nabla\phi(x^{+}),y-x\rangle$. Thus the entire expression simplifies to
$\langle\nabla\phi(x^{+})-\nabla\phi(x),\,y-x\rangle\;=\;\langle\nabla\phi(x)-\nabla\phi(x^{+}),\,x-y\rangle,$
which proves the desired identity. ∎
Definition 2 (Bregman projection).
Let $X\subset\mathcal{D}$ be closed and convex. For $y\in\mathcal{D}$, the (right) Bregman projection of $y$ onto $X$ is
$\Pi_{X}^{\phi}(y)\;\in\;\arg\min_{x\in X}D_{\phi}(x,y).$
Lemma 7 (Bregman Pythagorean theorem).
Let $X\subset\mathcal{D}$ be nonempty, closed, and convex, and let $x^{+}=\Pi_{X}^{\phi}(y)$. Then for all $x\in X$,
$D_{\phi}(x,y)\;\geq\;D_{\phi}(x,x^{+})\;+\;D_{\phi}(x^{+},y).$
Proof.
By first-order optimality for the convex problem $\min_{x\in X}D_{\phi}(x,y)$,
$\big\langle\nabla_{x}D_{\phi}(x,y)\big|_{x=x^{+}},\,x-x^{+}\big\rangle\;\geq\;0\quad\text{for all }x\in X.$
Since $\nabla_{x}D_{\phi}(x,y)=\nabla\phi(x)-\nabla\phi(y)$, we have
$\big\langle\nabla\phi(x^{+})-\nabla\phi(y),\,x-x^{+}\big\rangle\;\geq\;0.$
Apply Lemma 6 with $(x,y,z)=(x^{+},y,x)$ to obtain
$\big\langle\nabla\phi(x^{+})-\nabla\phi(y),\,x-x^{+}\big\rangle=D_{\phi}(x,x^{+})-D_{\phi}(x,y)+D_{\phi}(x^{+},y).$
Rearranging yields the stated inequality. ∎
Proof of Lemma 2
By construction, $(M_{T},\mathcal{F}_{T})$ is a mean-zero matrix-valued martingale:
$\mathbb{E}[D_{t}\mid\mathcal{F}_{t-1}]=0,\qquad\mathbb{E}[M_{T}\mid\mathcal{F}_{t-1}]=M_{t-1}.$
Because $\|c(x,a)\|$ is uniformly bounded,
$\|Y_{t}\|_{F}^{2}\leq d^{2}\,\sup_{x,a}\|c(x,a)\|^{4}<\infty.$
Hence $\mathbb{E}[\|D_{t}\|_{F}^{2}\mid\mathcal{F}_{t-1}]\leq C$ for some constant $C<\infty$, and therefore
$\mathbb{E}\|M_{T}\|_{F}^{2}=\sum_{t=1}^{T}\mathbb{E}\|D_{t}\|_{F}^{2}\leq CT.$
By orthogonality of martingale differences,
$\mathbb{E}\|M_{T}\|_{F}^{2}=\sum_{t=1}^{T}\mathbb{E}\|D_{t}\|_{F}^{2}=\sum_{t=1}^{T}\mathbb{E}\big[\mathbb{E}[\|D_{t}\|_{F}^{2}\mid\mathcal{F}_{t-1}]\big]\leq CT.$
Thus
$\mathbb{E}\left\|\frac{M_{T}}{T}\right\|_{F}^{2}=\frac{1}{T^{2}}\mathbb{E}\|M_{T}\|_{F}^{2}\leq\frac{C}{T}\;\to\;0,$ (52)
which shows $M_{T}/T\to 0$ in $L^{2}$ and hence in probability. Note that in the above equation we have considered the Frobenius norm. However, as all finite dimensional norms are equivalent, $M_{T}/T\xrightarrow{\mathbb{P}}0$ with respect to the operator norm as well. From our assumption it follows that there exists non-random weight vector $w^{\star}_{T}$ such that
$\frac{1}{T}\sum_{t=1}^{T}w_{t,k}-w^{\star}_{T,k}\;\stackrel{{\scriptstyle p}}{{\rightarrow}}\;0,$ (53)
for all experts $i\in[K]$. Now we observe that,
$\frac{B_{T}}{T}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}\mid\mathcal{F}_{t-1}]=\sum^{K}_{k=1}\left(\frac{1}{T}\sum_{t=1}^{T}w_{t,k}\right)\Sigma_{k}$
As all the elements of $\Sigma_{k}$ are uniformly bounded, we have the following string of inequalities,
$\left\|\frac{B_{T}}{T}-\frac{1}{T}{\Sigma^{\star}_{T}}\right\|_{op}$ $=\left\|\sum^{K}_{k=1}\left(\frac{1}{T}\sum_{t=1}^{T}w_{t,k}-w^{\star}_{T,k}\right)\Sigma_{k}\right\|_{op}$
$\leq\sum^{K}_{k=1}\bigg|\frac{1}{T}\sum_{t=1}^{T}w_{t,k}-w^{\star}_{T,k}\bigg|\left\|\Sigma_{k}\right\|_{op}$
$\lesssim\sum^{K}_{k=1}\bigg|\frac{1}{T}\sum_{t=1}^{T}w_{t,k}-w^{\star}_{T,k}\bigg|=\|\bar{w}_{T}-w^{\star}_{T}\|_{1}$ (54)
Therefore, our proof is complete.
Proof of Lemma 4
Recall that
$\widehat{g}_{t,k}=\,\ell_{t}\,\frac{\pi_{k}(a_{t}\mid x_{t})}{Q_{t}(a_{t}\mid x_{t})},\qquad Q_{t}(a\mid x_{t})=\sum_{j=1}^{K}w_{t,j}\,\pi_{j}(a\mid x_{t}),$
and assume $|\ell_{t}|\leq 1$ almost surely. By definition of the local dual norm,
$\|\widehat{g}_{t}\|_{w_{t},*}^{2}=\sum_{k=1}^{K}w_{t,k}\widehat{g}_{t,k}^{2}=\ell_{t}^{2}\sum_{k=1}^{K}w_{t,k}\,\frac{\pi_{k}^{2}(a_{t}\mid x_{t})}{Q_{t}^{2}(a_{t}\mid x_{t})}.$
Conditioning on $x_{t}$ and summing over $a_{t}\sim Q_{t}(\cdot\mid x_{t})$ gives
$\mathbb{E}\big[\|\widehat{g}_{t}\|_{w_{t},*}^{2}\mid x_{t}\big]$ $=\sum_{a\in\mathcal{A}}Q_{t}(a\mid x_{t})\,\Big\{\ell_{t}^{2}\sum_{k=1}^{K}w_{t,k}\,\frac{\pi_{k}^{2}(a\mid x_{t})}{Q_{t}^{2}(a\mid x_{t})}\Big\}$
$=\sum_{a\in\mathcal{A}}\ell_{t}^{2}\sum_{k=1}^{K}w_{t,k}\,\frac{\pi_{k}^{2}(a\mid x_{t})}{Q_{t}(a\mid x_{t})}$
$\leq\sum_{a\in\mathcal{A}}\ell_{t}^{2}\sum_{k=1}^{K}w_{t,k}\,\frac{\pi_{k}(a\mid x_{t})\,\max_{j}\pi_{j}(a\mid x_{t})}{Q_{t}(a\mid x_{t})}$
$=\sum_{a\in\mathcal{A}}\ell_{t}^{2}\,\max_{j}\pi_{j}(a\mid x_{t})\frac{\sum_{k=1}^{K}w_{t,k}\pi_{k}(a\mid x_{t})}{Q_{t}(a\mid x_{t})}$
$=\sum_{a\in\mathcal{A}}\ell_{t}^{2}\,\max_{j}\pi_{j}(a\mid x_{t})\;\leq\;|\mathcal{A}|\,\max_{x,a}\ell_{t}^{2}\;\leq\;|\mathcal{A}|,$
where we used $Q_{t}(a\mid x_{t})=\sum_{k}w_{t,k}\pi_{k}(a\mid x_{t})$ in the third line, and $\max_{x,a}\ell_{t}^{2}\leq 1$ in the last inequality. This completes the proof.